A hollow steel sphere ,weighing 200kg is floating on water.A weight 10 kg is to be placed on it in order to submerge when the temperature is 20°C. How much less weight is to be placed when temperature increases to 25°C?


γwater= 1.5 * 10-4            αsteel=1 * 10-5

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear gaurav

γwater= 1.5 * 10-4           

αsteel=1 * 10-5

γsteel = 3 * 10-5

let initial volume of sphere is Vo

and initial density of water in do

so initiall Vo volume of water replaced by sphere

so Vo do g = (200 + 10)g

    Vo do   = 210 ....................1

when temperature incresed

volume of sphere changes to V =Vo(1+γsteel *5)

density of water changes  to d = do(1-γwater *5)

let now w weight is placed

Vdg = (200 +w)g

Vd =200+w

Vo(1+γsteel *5)  X do(1-γwater *5)  =200 + w

210 (1+γsteel *5)  (1-γwater *5)  = 200 +w

210 (1+1.5 * 10-4(1-7.5 *10-4)  = 200 +w

  210( 1 -6 *10-4) = 200 +w

 10 -.1260 = w

w =9.874 kg

 difference in weight = 10 -w

                              = .1260 kg

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin